OpenStudy (anonymous):

If the area of a certain triangle is 38 square centimeters and the base is 8 centimeters, what is the height? Please show all of your work. A=1/2bh

OpenStudy (anonymous):

A=1/2bh. So when you plug in everything you get 38=1/2(8 x h). If you divide 38 in half to get rid of 1/2 you end up with 19= 8 x h. Divide 19 by 8 to get h and you end up with h= 9.5.